### v2

# Accelerated ADMM: Automated Parameter Tuning and Improved Linear Convergence

Abstract
This work studies the linear convergence of an accelerated scheme of the Alternating Direction Method of Multipliers (ADMM) for strongly convex and Lipschitz-smooth problems. We use the methodology of expressing the accelerated ADMM as a Lur’e system, i.e., an interconnection of a linear dynamical system in feedback with a slope-restricted operator, and we use Integral Quadratic Constraints to establish linear convergence. In addition, we propose several parameter tuning heuristics and their impact on the convergence rate through numerical analyses. Our new bounds show improved linear convergence rates compared to the vanilla algorithm and previous proposed accelerated variants, which is also empirically validated on a LASSO regression benchmark.
keywords:
ADMM, First-order optimization, Lur’e systems, Integral Quadratic Constraints
thanks: F. Jakob acknowledges the support of the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
1 Introduction
The Alternating Direction Method of Multipliers (ADMM) (Glowinski1975; boydAdmm) has emerged as a benchmark algorithm among first-order methods for solving convex composite optimization problems. Its ability to decompose the original problem into subproblems that can be solved in parallel made it especially popular in computation-heavy domains as signal processing, machine learning or distributed control.
It is long established that ADMM enjoys provable convergence guarantees in the convex setting. DengYin showed that for general convex problems ADMM achieves a $\mathcal{O}(1/k)$ convergence rate, while for the strongly convex and smooth case it attains a linear convergence rate, with optimally tuned parameters given in pontus_boyd. An alternative proof of linear convergence was presented by NishiharaADMM, who applied the Integral Quadratic Constraints (IQC) framework (Megretski1997; LessardRechtPackard; SchererEbenbauer) to numerically derive convergence rates and tune parameters. The proof frames the optimization algorithm as a Lur’e system (a feedback interconnection of an LTI system and monotone operators), and casts the convergence analysis as a stability problem. This framework offers the main advantage that also the analysis of altered schemes, such as over-relaxed variants, can be seamlessly carried out in an automated fashion as long as they can be framed in the Lur’e setting.
Also accelerated variants of ADMM (A-ADMM) have been proposed and analyzed in recent years. FastADMM showed that Nesterov acceleration leads to a convergence rate of $\mathcal{O}(1/k^{2})$ in the strongly convex settings. Sterck provided linear convergence guarantees when Anderson acceleration is applied. PatrinosStellaBemporad; PejcicJones proved linear convergence of accelerated Douglas–Rachford splitting and ADMM for strongly convex quadratic problems, showing that the optimal rate matches that of Nesterov’s fast gradient method. However, an analysis of the convergence rates of A-ADMM for general strongly convex and smooth problems is still unaddressed, as well as the question of whether other known rates from classical accelerated gradient methods can be matched.
Table 1: Contextualization of the results. Here, $\rho_{1}=1-\frac{1}{\sqrt{\kappa}}$ and $\rho_{2}=\sqrt{1-\frac{\sqrt{2\kappa-1}}{\kappa}}$with $\kappa=\frac{L}{m}$. NM and TM stand for Nesterov and Triple Momentum parameter selections (see Table 2).
Setting Objective $f$ Algorithm Rate $\rho$ Reference
$\min f(x)$ Quadr. Nesterov Method $\rho_{1}$ Nesterov2004
$\min f(x)$ $S(m,L)$ Nesterov Method $\rho_{2}$ safavi
$\min f(x)$ $S(m,L)$ Triple Momentum $\rho_{1}$ Freeman2028
(1) Quadr. A-ADMM (NM) $\rho_{1}$ PatrinosStellaBemporad; PejcicJones
(1) $S(m,L)$ A-ADMM (NM) Fig. 1(a) This work
(1) $S(m,L)$ A-ADMM (TM) Fig. 1(b) This work
In this work, we extend the IQC-based analysis of vanilla ADMM (NishiharaADMM) to the accelerated case, when one of the objectives is strongly convex and smooth. By casting the A-ADMM algorithm as a Lur’e system, we propose a semi-definite program (SDP) that provides a numerical tool to theoretically verify worst-case convergence rates of accelerated schemes. Using dynamic O’Shea-Zames-Falb (OZF) IQCs (BoczarLessardRecht), we propose a systematic parameter tuning procedure and demonstrate that Nesterov- and Triple-Momentum-inspired parameter selections attain rates close to those of accelerated first-order methods (cf. Table 1). Finally, we identify through grid search a new A-ADMM configuration with the fastest certified convergence rate, outperforming all vanilla ADMM methods. Our theoretical results are validated on a LASSO regression case study, demonstrating consistently improved convergence speed of our schemes compared to existing benchmark algorithms.
The remainder of the paper is organized as follows. Section 2 presents the problem setup, the ADMM formulation, and the required IQC preliminaries. Section 3 provides a Lur’e representation of the A-ADMM scheme. In Section 4, we study the convergence rates of this A-ADMM scheme under different parameter configurations. In Section 5, we provide a case study on LASSO regression. Finally, Section 6 concludes the paper.
Notation. The identity matrix of dimension $p$ is denoted as $I_{p}$. The gradient and subdifferential of a function $f$ are denoted by $\nabla f$ and $\partial f$, respectively. For $0<m\leq L<\infty$, we let $S_{p}(m,L)$ denote the class functions $f:\mathbb{R}^{p}\to\mathbb{R}$ that are $m$-strongly convex and have $L$-Lipschitz continuous gradients. The special case $S_{p}(0,\infty)$ corresponds to the set of proper, closed and convex functions. For a matrix $M$, its condition number is defined as $\kappa_{M}=\bar{\sigma}(M)/\underline{\sigma}(M)$, where $\bar{\sigma}(M)$ and $\underline{\sigma}(M)$ denote the largest and smallest singular values of $M$, respectively. A discrete-time LTI system maps the signals $u\mapsto y$ via the recursion $\xi_{k+1}=A\xi_{k}+Bu_{k},\;y_{k}=C\xi_{k}+Du_{k}$ as an $\xi_{0}$ dependent mapping, and will be compactly expressed as $y=Gu$ with $G={\small\left[\begin{array}[]{c|c}A&B\\[-1.0pt] \hline\cr C&D\end{array}\right]}$. We write $G\otimes I_{d}$ to define a lifted system $\left[\begin{array}[]{c|c}A\!\otimes\!I_{d}&B\!\otimes\!I_{d}\\[-1.0pt] \hline\cr C\!\otimes\!I_{d}&D\!\otimes\!I_{d}\end{array}\right]$. The series interconnection between two LTI systems $G_{1}:u\mapsto y$ and $G_{2}:y\mapsto w$ will be denoted as $G_{1}\cdot G_{2}:u\mapsto w$. The forward shift operator is denoted as $\mathbf{z}$.
2 PRELIMINARIES
2.1 Accelerated ADMM
We consider constrained convex optimization problems of the form
$\min_{x\in\mathbb{R}^{p},\;z\in\mathbb{R}^{q}}\;f(x)+g(z)\quad\text{s.t.}\quad Ax+Bz=c,$ (1)
where $f:\mathbb{R}^{p}\to\mathbb{R}$, $f\in S_{p}(m,L)$ and $g:\mathbb{R}^{q}\to\mathbb{R}$, $g\in S_{q}(0,\infty)$. The matrices $A\in\mathbb{R}^{p\times p}$, $B\in\mathbb{R}^{p\times q}$, and $c\in\mathbb{R}^{p}$ define linear equality constraints coupling the variables $x$ and $z$, where we particularly assume that $A$ is invertible and that $B$ has full column rank. Having a strongly convex and smooth component in (1) is common in many practical applications (DengYin), while restricting $A$ to be square and invertible is slightly restrictive, but also done in NishiharaADMM. We note that still many practical problem instances satisfy this structure, including consensus problems and various distributed optimization tasks (NotarstefanoNotarnicolaCamisa).
To address problem (1), we use the over-relaxed ADMM algorithm (boydAdmm)
$x_{k+1}$ $=\arg\min_{x}\;f(x)+\tfrac{1}{2\nu_{1}}\,\|Ax+B\hat{z}_{k}-c+\hat{\lambda}_{k}\|^{2}$ (2a)
$z_{k+1}$ $=\arg\min_{z}\;g(z)+\tfrac{1}{2\nu_{1}}\,\|\alpha Ax_{k+1}-(1-\alpha)B\hat{z}_{k}$ (2b)
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+Bz-\alpha c+\hat{\lambda}_{k}\|^{2}$
$\lambda_{k+1}$ $=\alpha Ax_{k+1}-(1-\alpha)B\hat{z}_{k}+Bz_{k+1}-\alpha c+\hat{\lambda}_{k}$ (2c)
as starting point, and augment it with a momentum-term
$\hat{z}_{k}$ $=z_{k}+\nu_{2}(z_{k}-z_{k-1})$ (3a)
$\hat{\lambda}_{k}$ $=\lambda_{k}+\nu_{2}(\lambda_{k}-\lambda_{k-1}),$ (3b)
as in FastADMM; PejcicJones. Here, $\nu_{1}>0$ is a step size, $\nu_{2}\geq 0$ a momentum parameter and $\alpha\geq 0$ the so-called over-relaxation parameter. We recover the nominal A-ADMM by setting $\alpha=1$, and the vanilla, non-accelerated ADMM with $\nu_{2}=0$. Both $\alpha$ and $\nu_{2}$ have been shown to improve the convergence speed for suitable choices (Eckstein1992; PatrinosStellaBemporad). The over-relaxation parameter $\alpha$ is chosen in the interval $(0,2]$, while $\nu_{2}$ is often selected according to Nesterov’s schemes (Nesterov2004), e.g., $\nu_{2}=(\sqrt{L}-\sqrt{m})/(\sqrt{L}+\sqrt{m})$ (PatrinosStellaBemporad), or as an iteration dependent quantity (FastADMM). To explore alternative parameter selection strategies we employ the IQC framework.
2.2 Integral Quadratic Constraints for Convex Functions
Algorithm (2)-(3) will be analyzed as a dynamical system converging to points satisfying the first-order optimality conditions. We will make use of the well-known fact that (sub)gradients of (strongly) convex functions are monotone and, thus, their input/output relation satisfy IQCs for slope-restricted operators (LessardRechtPackard). To be more precise, they have been shown to fulfill so-called $\rho$-hard O’Shea-Zames–Falb IQCs.
Proposition 1 ($\rho$-OZF IQC, BoczarLessardRecht)
Let $f\in S_{p}(m,L)$, $g\in S_{p}(0,\infty)$ and let $\rho\in(0,1)$ be an exponential discount factor. For $n_{\mathrm{OZF}}\in\mathbb{N}_{0}$, let $\{h^{f}_{\tau}\}_{\tau=0}^{n_{\mathrm{OZF}}}$ and $\{h^{g}_{\tau}\}_{\tau=0}^{n_{\mathrm{OZF}}}$ be sequences of filter coefficients satisfying
$h_{\tau}^{*}\leq 0\,\,\,\forall\tau\neq 0,\quad\sum_{\tau=0}^{n_{\mathrm{OZF}}}\rho^{-2\tau}h_{\tau}^{*}\geq 0,\quad*\in\{f,g\}$ (4)
and define the parametrized LTI filters
$\Psi_{g}(h^{g})=\begin{bmatrix}{\tiny\displaystyle\sum_{\tau=0}^{n_{\mathrm{OZF}}}}h_{\tau}^{g}\mathbf{z}^{-\tau}&0\\ 0&1\end{bmatrix},\,\Psi_{f}(h^{f})=\Psi_{g}(h^{f})\cdot\begin{bmatrix}L&-1\\ -m&1\end{bmatrix}$. (5)
Let $\{a_{k}\},\{b_{k}\}$ be some $p$-dim. square summable sequences and let $a^{\star},b^{\star}\in\mathbb{R}^{p}$ be constant references. For $\gamma_{k}\in\partial g(b_{k})$ and $\gamma^{\star}\in\partial g(b^{\star})$, define
$\tilde{a}_{k}$ $=a_{k}-a^{\star},$ ${\nabla\tilde{f}}_{k}$ $=\nabla f(a_{k})-\nabla f(a^{\star}),$
$\tilde{b}_{k}$ $=b_{k}-b^{\star},$ $\tilde{\gamma}_{k}$ $=\gamma_{k}-\gamma^{\star}.$
Finally, define the filtered sequences
$\psi_{1,k}=(\Psi_{f}\otimes I_{p})\begin{bmatrix}\tilde{a}_{k}\\ {\nabla\tilde{f}}_{k}\end{bmatrix},\quad\psi_{2,k}=(\Psi_{g}\otimes I_{p})\begin{bmatrix}\tilde{b}_{k}\\ \tilde{\gamma}_{k}\end{bmatrix}$
and $M=\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]$. Then, for all $T\geq 0$, it holds
$\begin{aligned} &\sum_{k=0}^{T}\rho^{-2k}\psi_{i,k}^{\top}(M\otimes I_{p})\psi_{i,k}\geq 0,\quad i=1,2.\end{aligned}$ (6)
The filter coefficients $h^{f}$, $h^{g}$ will serve as degree-of-freedom in the resulting SDP, subject to the convex constraint (4). In theory, the filter dimension $n_{\mathrm{OZF}}$ can be infinite, but is chosen finite in practice to yield finite-dimensional state-space realizations. A choice of $n_{\mathrm{OZF}}=0$ leads to static filters $\Psi_{f},\Psi_{g}$ and pointwise satisfaction of the inequalities (6), which in literature is often referred to as sector IQC (LessardRechtPackard).
2.3 Stability Analysis of First-Order Algorithms
Consider the following Lur’e representation of a first-order algorithm
$\begin{aligned} \xi_{k+1}&=(\hat{A}\otimes I_{p})\xi_{k}+(\hat{B}\otimes I_{p})w_{k}\\ \hat{v}_{k}&=(\hat{C}\otimes I_{p})\xi_{k}+(\hat{D}\otimes I_{p})w_{k}\end{aligned}$ (7a)
with state $\xi_{k}\in\mathbb{R}^{n_{\xi}}$, output $\hat{v}_{k}=v_{k}+\bar{v}$ for some constant offset $\bar{v}$, and
$v_{k}=\begin{bmatrix}v_{1,k}\\ v_{2,k}\end{bmatrix},\quad w_{k}=\begin{bmatrix}\nabla\hat{f}(v_{1,k})\\ \gamma_{k}\end{bmatrix},\,\gamma_{k}\in\partial\hat{g}(v_{2,k}),$ (7b)
[…]
$\tilde{v}=\Biggl(\underbrace{\left[\begin{array}[]{c|c}\hat{A}&\hat{B}\\ \hline\cr\hat{C}&\hat{D}\end{array}\right]}_{=:G}\otimes I_{p}\Biggr)\tilde{w}$. (8)
[…]
$\psi_{k}\triangleq\begin{bmatrix}\psi_{1,k}\\ \psi_{2,k}\end{bmatrix}=(\Psi\otimes I_{p})\begin{bmatrix}\tilde{v}_{k}\\ \tilde{w}_{k}\end{bmatrix}$, (9)
with $\psi_{1,k},\psi_{2,k}$ as in Proposition 1 (with $a\triangleq v_{1},\,b\triangleq v_{2}$). Define the augmented plant as the mapping $\tilde{w}\mapsto\psi$, realized as
$\Psi\cdot\begin{bmatrix}G\\ I_{2}\end{bmatrix}\triangleq\left[\begin{array}[]{c|c}\mathbf{A}&\mathbf{B}\\ \hline\cr\mathbf{C}&\mathbf{D}\end{array}\right].$ (10)
Then the following convergence theorem holds.
Theorem 2
Take $\rho\in(0,1)$, and filters $\Psi_{1},\Psi_{2}$ that satisfy Proposition 1 with $m=\hat{m},L=\hat{L}$. Form the augmented plant (10). If there exist $P\succ 0$ and filter coefficients of $\Psi_{1},\Psi_{2}$ such that the matrix inequality
$\begin{bmatrix}\mathbf{A}^{\top}P\mathbf{A}-\rho^{2}P&\mathbf{A}^{\top}P\mathbf{B}\\ \mathbf{B}^{\top}P\mathbf{A}&\mathbf{B}^{\top}P\mathbf{B}\end{bmatrix}+\begin{bmatrix}\mathbf{C}^{\top}\\ \mathbf{D}^{\top}\end{bmatrix}\begin{bmatrix}M&0\\ 0&M\end{bmatrix}\begin{bmatrix}\mathbf{C}&\mathbf{D}\end{bmatrix}\;\preceq\;0$ (11)
holds, then (7) is exponentially stable with rate $\rho$, i.e.,
$\|\xi_{k}-\xi^{\star}\|\leq\sqrt{\kappa_{P}}\,\rho^{k}\|\xi_{0}-\xi^{\star}\|.$ (12)
The proof comes as a straightforward extension of (LessardRechtPackard, Theorem. 4) to the two-operator case, which particularly exploits the positivity constraint (6). Eq. (11) is independent of the problem dimension $p$, and can be used to determine the minimum worst-case convergence rate $\rho$ via bisection. Note that Theorem 2 recovers (NishiharaADMM, Theorem. 6) for $n_{\mathrm{OZF}}=0$.
3 A-ADMM as a Dynamical System
Building on the previous derivations, we now formulate the A-ADMM (2)-(3) as a Lur’e system (7). In line with NishiharaADMM, we introduce the coordinate change $r_{k}=Ax_{k}$, $s_{k}=Bz_{k}$. Moreover, define
$\hat{f}=f\circ A^{-1},\quad\hat{g}=g\circ B^{\dagger}+\mathcal{I}_{\mathrm{im}\,B},$ (13)
[…]
with
$\hat{m}=\tfrac{m}{\bar{\sigma}^{2}(A)},\quad\hat{L}=\tfrac{L}{\underline{\sigma}^{2}(A)}.$ (14)
Accordingly, we define the condition number of the problem as $\kappa=\frac{\hat{L}}{\hat{m}}=\kappa_{f}\kappa_{A}^{2}$.
Consequently, the updates (2a),(2b),(3a) can be rewritten as
$x_{k+1}$ $=A^{-1}\arg\min_{r}\;\hat{f}(r)+\tfrac{1}{2\nu_{1}}\,\lVert r+\hat{s}_{k}-c+\hat{\lambda}_{k}\rVert^{2}$ (15a)
$z_{k+1}$ $=B^{\dagger}\arg\min_{s}\;\hat{g}(s)+\tfrac{1}{2\nu_{1}}\,\lVert\alpha r_{k+1}$- (15b)
$\hskip 99.58464pt(1-\alpha)\hat{s}_{k}+s-\alpha c+\hat{\lambda}_{k}\rVert^{2}$
$\hat{s}_{k}$ $=s_{k}+\nu_{2}(s_{k}-s_{k-1}).$ (15c)
Using the definition of the proximal operator
$\mathrm{prox}_{f}(z):=\arg\min_{x}f(x)+\frac{1}{2}\|x-z\|^{2},$
we can summarize the A-ADMM in transformed coordinates compactly as
$r_{k+1}$ $=\operatorname{prox}_{\nu_{1}\hat{f}}\!\big(c-\hat{s}_{k}-\hat{\lambda}_{k}\big)$ (16a)
$s_{k+1}$ $=\operatorname{prox}_{\nu_{1}\hat{g}}\!\big(\alpha c-\alpha r_{k+1}-(\alpha-1)\hat{s}_{k}-\hat{\lambda}_{k}\big)$ (16b)
$\lambda_{k+1}$ $=\alpha r_{k+1}+(\alpha-1)\hat{s}_{k}+s_{k+1}-\alpha c+\hat{\lambda}_{k},$ (16c)
with $\hat{s}_{k},\hat{\lambda}_{k}$ as in (15c), (3b). Next, we bring the recursion (16) into the form (7).
Proposition 3
Define the state $\xi_{k}\triangleq\begin{bmatrix}\lambda_{k-1}^{\top}&s_{k-1}^{\top}&\lambda_{k}^{\top}&s_{k}^{\top}\end{bmatrix}^{\top}$, output $v_{k}\triangleq\begin{bmatrix}r_{k+1}^{\top}&s_{k+1}^{\top}\end{bmatrix}^{\top}$, offset $\bar{v}\triangleq\begin{bmatrix}-c^{\top}&0\end{bmatrix}^{\top}$, and input $w_{k}\triangleq\begin{bmatrix}\nabla\hat{f}(r_{k+1})^{\top}&\gamma_{k}^{\top}\end{bmatrix}^{\top}$ for any $\gamma_{k}\in\partial\hat{g}(s_{k+1})$. Then the sequences $\xi_{k}$, $w_{k}$, and $\hat{v}_{k}=v_{k}+\bar{v}$ satisfy (7) with the matrices
[…]
{pf}
Apply the first-order optimality condition to the $r_{k+1}$ and $s_{k+1}$ updates (16a)-(16b)
$0$ $=\nu_{1}\nabla\hat{f}(r_{k+1})+r_{k+1}-c+\hat{s}_{k}+\hat{\lambda}_{k},$ (18a)
$0$ $\in\nu_{1}\partial\hat{g}(s_{k+1})+s_{k+1}-\alpha c+\alpha r_{k+1}+(\alpha-1)\hat{s}_{k}+\hat{\lambda}_{k}.$ (18b)
Take some $\gamma_{k}\in\partial g(s_{k+1})$, solve for $r_{k+1}$ and $s_{k+1}$, and plug it into the dual update $\lambda_{k+1}$, to get
$r_{k+1}$ $=-\hat{s}_{k}-\hat{\lambda}_{k}+c-\nu_{1}\nabla\hat{f}(r_{k+1}),$ (19a)
$s_{k+1}$ $=\hat{s}_{k}+(\alpha-1)\hat{\lambda}_{k}+\alpha\nu_{1}\nabla\hat{f}(r_{k+1})-\nu_{1}\gamma_{k},$ (19b)
$\lambda_{k+1}$ $=-\nu_{1}\gamma_{k}.$ (19c)
Now plug in the definitions of $\hat{s}_{k},\hat{\lambda}_{k}$ and observe that $s_{k+1},\lambda_{k+1}$ and $r_{k+1}-c$ can all be written as linear combinations of the state $\xi_{k}$ and input $w_{k}$. It is then straightforward to bring $\xi_{k+1}$ and $\hat{v}_{k}$ into the matrix form (7). ∎
Remark 4
In contrast to NishiharaADMM, we deliberately retain the ADMM step-size $\nu_{1}$ in the state-space description, instead of putting it in the sector bounds. The momentum parameter $\nu_{2}$ becomes an additional degree-of-freedom.
4 Parameter Selection And Additional Modifications
In unconstrained first-order methods, accelerated algorithms are parametrized by a step-size, a damping-parameter, and an extrapolation-parameter, where each quantity has provable optimal tunings (Freeman2028). For ADMM in contrast, convergence proofs and parameter selection of $\nu_{1},\nu_{2},\alpha$ have been traditionally difficult. In this section we explore different parameter selection heuristics and show the influence on the convergence rates $\rho$ that are obtained from Theorem 2. As initial heuristics, we adopt the same parameter tuning schemes of accelerated gradient methods. Next, we perform a grid-search to identify potentially superior configurations.
4.1 Nesterov and Triple Momentum Parameters
We investigate the use of the parameterization of the Nesterov method (NM) (Nesterov2004) and the Triple Momentum algorithm (TM) (Freeman2028), which is known to attain the fastest possible convergence rate for accelerated gradient methods. The parameter selections are summarized in Table 2. We note that it had already been proposed to tune A-ADMM based on the NM parameter set (PejcicJones). We plot the resulting $\rho$ obtained by Theorem 2 as a function of the condition number $\kappa=\hat{L}/\hat{m}$. We increase the O’Shea-Zames-Falb order until $\rho$ does not improve anymore, giving $n_{\mathrm{OZF}}=6$ in all experiments [^1] Code for all experiments can be found on https://github.com/col-tasas/2025-accelerated-admm .
For the NM parameter set, observe in Fig. 1(a) that the convergence rates indeed approach the optimal rate of Nesterov’s accelerated method, given by $\rho={\scriptsize\sqrt{1-\frac{\sqrt{2\kappa-1}}{\kappa}}}$ (safavi). For the TM parameter set, observe in Fig. 1(b) that the convergence rates also approach the optimal rate of Triple Momentum, given by $\rho=1-\frac{1}{\sqrt{\kappa}}$, However, from $\kappa\approx 18$ onwards, the rates degrade and eventually become larger than one. Thus, no statement about the convergence for higher condition ratios can be done. We now propose a modification to the ADMM update to alleviate this issue.
Table 2: Parameter selection for A-ADMM as a function of $\hat{m},\hat{L}$, with $\rho:=1-1/\sqrt{\kappa}$.
Method $\nu_{1}$ $\nu_{2}$
NM $\frac{1}{\hat{L}}$ $\frac{\sqrt{\hat{L}}-\sqrt{\hat{m}}}{\sqrt{\hat{L}}+\sqrt{\hat{m}}}$
TM $\frac{1+\rho}{\hat{L}}$ $\frac{\rho^{2}}{2-\rho}$
[…]
(a) Using NM parameters.
[…]
(b) Using TM parameters.
Figure 1: Comparison of optimal $\rho$ values for NM and TM parameterizations.
4.2 $\lambda$-Damping
Observe in (19c) that, unlike the $s_{k}$-dynamics, the $\lambda_{k}$ dynamics depends solely on the subdifferential of $\hat{g}$ and is therefore undamped. We propose to introduce an additional damping of the dual variable $\lambda_{k}$ as
$\lambda_{k+1}=d\lambda_{k}-d\nu_{1}\gamma_{k},$
with some damping parameter $d\in(0,1)$. From the point of view of the original algorithm iterations we can interpret this modifications by replacing the dual update in (2) with
$\lambda_{k+1}=d(\alpha Ax_{k+1}-(1-\alpha)Bz_{k}+Bz_{k+1}-\alpha c+\hat{\lambda}_{k}+\lambda_{k}).$ (20)
The state-space representation of Proposition 3 is modified accordingly, with the only change occurring in the $\hat{A}$ and $\hat{B}$ matrix. We note that the fixed-point of the $\lambda$-damped scheme still satisfies the optimality conditions of (1), as the damping vanishes at optimality.
The results of A-ADMM with TM parameters and additional $\lambda$-damping with $d=0.1$ are shown in Fig. 1(b). We observe that convergence can now be guaranteed for all condition ratios, and moreover, we observe analogous to the NM case that the convergence rate curve approaches the optimal rate of Triple Momentum, given by $\rho=1-\frac{1}{\sqrt{\kappa}}$. This suggests that the additional damping of the dual variable extends the range of provable convergence for TM A-ADMM, and that the strong connection between A-ADMM and accelerated first-order methods is also preserved in the strongly-convex and smooth setting.
4.3 Grid Search Results
We now perform a grid search to identify potentially improved configurations of $(\nu_{1},\nu_{2})$. Our search is set up as a $20\times 20$ grid centered around the TM parameters. As commonly done in strongly-convex and smooth optimization problems, we parametrize the parameters as a function of the condition ratio $\kappa$. For each $\kappa$ we select $(\nu_{1},\nu_{2})$ as the pair for which the minimal $\rho$ can be obtained with Theorem 2 via bisection. The procedure is repeated for the over-relaxed A-ADMM (OR-A-ADMM), with an additional grid over the relaxation parameter $\alpha$ in the range $(0,4]$. Note that even though the standard theory would prevent from exploring the range $(2,4)$, we can still perform this interval extension and check for linear convergence by the feasibility of (11).
[…]
(a) Results for $\nu_{1}$.
[…]
(b) Results for $\nu_{2}$.
Figure 2: Results of the grid search for ($\nu_{1}$, $\nu_{2}$) vs. TM parameters from Table 2.
Fig. 2 shows the grid search results for $\nu_{1}$ and $\nu_{2}$. We found that the stepsize parameter $\nu_{1}$ found via grid search exactly coincides with the TM value. In contrast, the step-size parameter $\nu_{2}$ exhibits a different behavior. The same holds for the over-relaxed case. Applying symbolic regression (pysr) reveals that $\nu_{2}$ can be approximated by
$\nu_{2}$ $\approx\left(\frac{\kappa+0.08}{\kappa+49.9}\right)^{\frac{1}{4}}-0.2$ (21)
$\nu_{2}^{\mathrm{OR}}$ $\approx\left(\frac{0.66}{\kappa+11.97}\right)\kappa+0.06,$ (22)
for the nominal and over-relaxed settings, respectively. The over-relaxation parameter $\alpha$ was found to be optimal for $\alpha\approx 1.45$ across all condition numbers.
Fig. 3 summarizes the convergence rates obtained from the grid search (GS) parameters, in comparison to the vanilla ADMM method and the accelerated versions with NM and TM parameters. By design, the A-ADMM with GS-parameters gives the best linear convergence rate of among all settings. Fig. 4 complements the results with the over-relaxed variants. We observe that the over-relaxed variants constantly outperform the non-over-relaxed ones; and that the accelerated variants constantly outperform the non-accelerated variants.
[…]
Figure 3: Comparison of optimal $\rho$ values for ADMM and A-ADMM with different parameter configurations.
[…]
Figure 4: Comparison of optimal $\rho$ values nominal and over-relaxed ADMM versions.
5 Case Study
We test the A-ADMM schemes designed based on the theoretical analyses in the previous section on the $\ell_{1}$-regularized least-squares problem, also known as LASSO Regression, and compare them with available algorithms (boydAdmm). We use the benchmark problem
$\min_{x,z}\quad$ $\tfrac{1}{2}\|Fx-b\|_{2}^{2}+\tau\|z\|_{1}$
s.t. $x=z,$ (23)
where $F\in\mathbb{R}^{500\times 100}$ is chosen as a tall and full column rank matrix with entries that are sampled i.i.d. from an isotropic Gaussian distribution $\mathcal{N}(0,1)$, with normalized columns. The vector $b$ is generated as $b=Fw^{0}+\varepsilon$, where $w^{0}\in\mathbb{R}^{100}$ is sparse, containing 50 non-zeros drawn from $\mathcal{N}(0,1)$ and $\varepsilon\sim\mathcal{N}(0,10^{-3}I_{500})$. The regularization parameter $\tau>0$, is set to $\tau=0.01$. Problem (23) fits the general problem (1), where the first term is $m$-strongly convex and $L$-smooth with $m=\underline{\sigma}^{2}(F^{\top}F)$ and $L=\bar{\sigma}^{2}(F^{\top}F)$.
Applying the accelerated over-relaxed ADMM results in
$x_{k+1}$ $=(\nu_{1}F^{\top}F+I)^{-1}\left(-\hat{z}_{k}-\hat{\lambda}_{k}+\nu_{1}F^{\top}b\right),$ (24a)
$z_{k+1}$ $=\mathcal{S}_{\tau\nu_{1}}\left(-\alpha\,x_{k+1}-(\alpha-1)\hat{z}_{k}-\hat{\lambda}_{k}\right),$ (24b)
$\lambda_{k+1}$ $=\alpha\,x_{k+1}+(1-\alpha)\,\hat{z}_{k}+z_{k+1}+\hat{\lambda}_{k},$ (24c)
$\hat{z}_{k}$ $=z_{k}+\nu_{2}\,(z_{k}-z_{k-1}),$ (24d)
$\hat{\lambda}_{k}$ $=\lambda_{k}+\nu_{2}\,(\lambda_{k}-\lambda_{k-1}).$ (24e)
with the soft-thresholding operator $\mathcal{S}_{\tau\nu_{1}}$ that is (element-wise) defined as
$\bigl[\mathcal{S}_{\tau\nu_{1}}(\mathbf{y})\bigr]_{i}=\operatorname{sign}(\mathbf{y}_{i})\,\max\bigl(|\mathbf{y}_{i}|-\tau\nu_{1},0\bigr).$ (25)
The convergence is assessed through the normalized iterate error
$\Delta_{k}:=\frac{\|x_{k}-x^{\star}\|_{2}}{\|x_{0}-x^{\star}\|_{2}},$ (26)
where $x^{\star}$ is the solution of (23).
In particular, we study the convergence speed of A-ADMM and OR-A-ADMM, both with TM and GS parameters, against the widely used LASSO benchmark solver FISTA (FISTA) as well as ADMM and OR-ADMM. The iterate error computed on one random instance of (23) is shown in Fig. 5, while Table 3 summarizes the average number of iterations required to reach $\Delta_{k}<10^{-5}$ over 200 random instances.
[…]
Figure 5: Empirical performance of different A-ADMM schemes and other benchmark algorithms applied to a random instance of LASSO regression.
Table 3: Average iterations to reach $\Delta_{k}<10^{-6}$ over 200 random LASSO instances.
Method Average iterations
ADMM 46.19
A-ADMM (TM) 46.24
FISTA 39.3
OR-ADMM 37.48
A-ADMM (GS) 36.66
OR-A-ADMM (GS) 34.98
Both Fig. 5 and Table 3 show that the performance of the A-ADMM methods differ profoundly based on their parameter selection. While a TM tuning is empirically showing no advantage over an optimally tuned vanilla ADMM, we see that a GS tuning gives a substantial advantage. In addition, we observe that the OR-ADMM with GS tuning consistently outperforms all other methods. Notably, the ordering of algorithms evaluated on the empirical performance and evaluated on the worst-case convergence rates stays the same, suggesting that the worst-case convergence rate is a suitable metric for algorithm comparison.
6 Conclusions
This paper used the IQC-based algorithm analysis framework to show linear convergence rates of different accelerated ADMM variants. We proposed different parameter tunings, and show that the additional momentum parameters allows to outperform vanilla ADMM. The fastest version both in terms of certified convergence rate and empirical performance was found using a grid search over the parameter space, giving the fastest version of A-ADMM in the strongly-convex-smooth setting to date. We moreover showed the close connection between A-ADMM parameters and accelerated gradient methods from unconstrained optimization. Future works could focus on theoretic derivations of the optimal A-ADMM parameters, and the development of a systematic and robust A-ADMM synthesis procedure.